OpenStudy (anonymous):

How would you translate the graph of y=-x squared to produce the graph of y=-(x-6) squared?

OpenStudy (anonymous):

Okay: move the graph of -x^2 over 6 units to the right.

OpenStudy (anonymous):

is there step by steps to solve it?

OpenStudy (anonymous):

Not really: you need to think that x is changed out for x-6 Therefore x1 = x2 -6 x2 = x1 +6: shift the graph 6 to the right

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

by x2 you mean x squared right?

OpenStudy (anonymous):

I mean \[x _{2}\]

OpenStudy (anonymous):

you mean x^2?